Meeting 2025_10_17
Jan, William, Chiara, Daniele, Stephane
TS2 access:
- Daniele made an access this morning and had an overview:
- Dust is accumulating à clean up to foresee for next YETS (mirrors, filters, etc…)
- Nothing to do so far
- Nothing urgent to buy à LS3 together with new BSRT hardware
- Access to take a few measurment for a cover for HSE to be happy 😊 (William/Stephane)
- Replacement of target support and target (Stephane/Daniele)
YETS access:
- Clean up optical elements (Daniele/Stephane)
- Install new cover (William)
- Visite HSE (C.Gaignant and O.Tyson) to finalise the Safety Assessment ! (to be launched by Stephane or William?)
- Visite with HSE (?) / SCE (C.Biot) to discuss on site about: Hutch and optical lines, etc… (to be launched by Stephane or William?)
- Discussion with BGI to make sure of definitive use of carrotage (all)
- Christelle confirmed to Chiara the 2m height aperture allocation !! Since Germana said if HSE is OK….
- BSRA
- David confirmed no BSRA specific actions (except standard maintenance/check at restart)
ECR new BSR:
- Priority 1 is vacuum part
- BSRTR: Actual design (flange reduced to 150mm) still needs to be adapted for integration à vacuum chamber itself à Ben + Stephane next week
-
- Compressed air
- BTV electronics for BSRTR CTRL à cabling on US45 to optimize available electronics (Stephane to update DIC)
- Rack request with all services (power, network, etc…) à inside or outside hutch?
- Layout hutch
Test jitter in position HV design push-pull
See slide (Stephane) --> jitter is max +/-1mm over 16m
There are minutes attached to this event.
Show them.